Ram Das of Hyderabad consigned goods costing Rs. 72,000 to Prakash of Cochin at a pro-forma invoice price which is cost plus a profit of 1/6th on invoice price. The consignor paid Rs. 1,800 as insurance and other charges. Prakash received the goods and paid Rs. 3,000 for freight and other charges. He was allowed 3% commission on gross sales. 3/4th of the goods were sold at 33.33% profit on cost, half of which were credit sales. Half of the balance was stolen, but the stock being insured, a claim was lodged for Rs. 8,000 and was settled for Rs. 7,000. Balance of stock was valued at proforma invoice price. Write up the Consignment and the Abnormal Loss Accounts.
Answer #1

Dr Consignment A/c Cr

Particulars Amount Particulars Amount
To goods sent on consignment 86400 By prakash A/c (sales) 72000
To Cash(insurance exp) 1800 By Abnormal loss A/c 9600
To Prakash A/c (freight) 3000
To Prakash A/c (commision) 2160
By Inventory on Consignment 11520
To inventory reserve(loading) 1920 By goods sent on consignment(loading) 14400
To profit and Loss A/c 12240
107520 107520

Dr Abnormal loss A/c cr

particulars Amount Particulars Amount
To Consignment A/c 9600 By insurance Company 7000
To profit and loss A/c 2600
9600 9600

Workings

goods sent on consignment=72000/5+72000=86400

goods sent on consignment(loading)=86400-72000=14400

Calculation of Closing inventory at invoice price

Inventory =(72000)*1/4-loss=18000-18000/2=9000

proportion expenses=(3000+1800)*9000/72000=600

Total closing stock at cost=9000+600=9600

Closing inventory at invoice price=9600+9600/5=11520

Inventory reserve loading=11520-9600=1920

Abnormal loss=9000+1/8(3000+1800)=9600